# EssayJudge: A Multi-Granular Benchmark for Assessing Automated Essay Scoring Capabilities of Multimodal Large Language Models

### v1

@@ Abstract
Automated Essay Scoring (AES) plays a crucial role in educational assessment by providing scalable and consistent evaluations of writing tasks. However, traditional AES systems face three major challenges: ❶ reliance on handcrafted features that limit generalizability, ❷ difficulty in capturing fine-grained traits like coherence and argumentation, and ❸ inability to handle multimodal contexts. In the era of Multimodal Large Language Models (MLLMs), we propose EssayJudge, the first multimodal benchmark to evaluate AES capabilities across lexical-, sentence-, and discourse-level traits. By leveraging MLLMs’ strengths in trait-specific scoring and multimodal context understanding, EssayJudge aims to offer precise, context-rich evaluations without manual feature engineering, addressing longstanding AES limitations. Our experiments with 18 representative MLLMs reveal gaps in AES performance compared to human evaluation, particularly in discourse-level traits, highlighting the need for further advancements in MLLM-based AES research. Our dataset and code will be available upon acceptance.
@@
Jiamin Su1,*, Yibo Yan1,3,*, Fangteng Fu1, Han Zhang1, Jingheng Ye4, Xiang Liu1,3, Jiahao Huo1, Huiyu Zhou2, Xuming Hu1,2,3, [^2] Corresponding author. 1The Hong Kong University of Science and Technology (Guangzhou) 2Guangxi Zhuang Autonomous Region Big Data Research Institute 3The Hong Kong University of Science and Technology 4Tsinghua University jsu360@connect.hkust-gz.edu.cn, yanyibo70@gmail.com, xuminghu@hkust-gz.edu.cn
@@ 2.3 Multimodal Large Language Models
MLLMs have brought significant advancements to diverse tasks and applications Xi et al. (2023); Huo et al. (2024); Yan et al. (2024d); Yan and Lee (2024); Zou et al. (2025); Dang et al. (2024). Proprietary MLLMs such as GPT-4o Hurst et al. (2024) and Gemini-1.5 DeepMind (2024b) have shown remarkable capabilities in multimodal challenges, excelling in areas such as multimodal reasoning and QA Chang et al. (2024); Yan et al. (2024c, b); Zheng et al. (2024); Yan et al. (2025). At the same time, Open-source MLLMs have made considerable strides. For instance, LLaVA-NEXT Liu et al. (2024) utilizes a pretrained vision encoder to generate visual embeddings, which are then aligned with text embeddings through a lightweight adapter, enabling effective multimodal understanding. Similarly, MLLMs such as Qwen2-VL Wang et al. (2024), DeepSeek-VL Lu et al. (2024a), InternVL Chen et al. (2024, 2025), MiniCPM Hu et al. (2024), Ovis Lu et al. (2024b), LLaMA3 Dubey et al. (2024) and Yi-VL Young et al. (2024) implement innovative projection techniques to combine visual and textual features effectively, enabling many multimodal applications. These models showcase the growing potential of MLLMs in advancing both research and practical applications that rely on multimodal data Qu et al. (2025); Zou et al. (2024); Zhou et al. (2024); Huang et al. (2024). Therefore, we introduce EssayJudge, a novel benchmark designed to evaluate MLLMs’ capability to score essays with multimodal context, paving the way for AGI systems Xiao et al. (2024b); Tate et al. (2024); Yan et al. (2024a).
@@ 3.1 Data Collection
3.2 Data Annotaion Scheme
@@ 4.1 Experimental Setup
We meticulously categorized diverse MLLMs into distinct groups to assess their capabilities in trait-specific AES. (i) The Open-Source MLLMs category encompassed models such as Yi-VL Young et al. (2024), Qwen2-VL Wang et al. (2024), DeepSeek-VL Lu et al. (2024a), LLaVA-NEXT Liu et al. (2024), InternVL2 Chen et al. (2024), InternVL2.5 Chen et al. (2025), MiniCPM-V2.6 Hu et al. (2024), MiniCPM-LLaMA3-V2.5 Hu et al. (2024), Ovis1.6-Gemma2 Lu et al. (2024b), and LLaMA-3.2-Vision Dubey et al. (2024), each demonstrating their unique strengths and capabilities in multi-granular essay scoring. (ii) The Closed-Source MLLMs featured proprietary models like Qwen-Max Team (2024), Step-1V StepFun (2024), Gemini-1.5-Pro DeepMind (2024b), Gemini-1.5-Flash DeepMind (2024a), Claude-3.5-Haiku Anthropic (2024a), Claude-3.5-Sonnet Anthropic (2024b), GPT-4o-mini OpenAI (2024), and GPT-4o Hurst et al. (2024), providing a comparison point for the performance of models that are not publicly accessible. (iii) Lastly, the Human Performance category served as a benchmark for human-level intelligence, enabling us to assess how closely MLLMs emulate human cognitive abilities (More details in Appendix C.2). The detailed prompts for MLLMs and sources of MLLMs are provided in Appendix C.3 and C.4.
@@ Appendix C Additional Experimental Details
C.1 Multimodal Ablation Study
We also evaluated GPT-4o-mini. Figure 12 shows that GPT-4o-mini exhibited a decline in nine out of ten traits when image information was removed, with the exception of lexical diversity. This finding further underscores the importance of multimodal inputs, as visual features provide essential evaluative dimensions that text-only approaches cannot capture, especially when images contain critical supporting information.
C.2 Human Performance Evaluation
@@
C.3 Prompt for MLLM Evaluation
@@ Appendix F More Multimodal Essay Scoring Examples
Table 7: Rubrics for evaluating the essay lenth of the essays.
@@
Table 11: Rubrics for evaluating the lexiacal accuracy of the essays.
@@
Table 12: Rubrics for evaluating the lexiacal diversity of the essays.
@@
Figure 12: GPT-4o-mini’s QWK values across traits for for text-only and multimodal settings.

### v2

@@ Abstract
Automated Essay Scoring (AES) plays a crucial role in educational assessment by providing scalable and consistent evaluations of writing tasks. However, traditional AES systems face three major challenges: ❶ reliance on handcrafted features that limit generalizability, ❷ difficulty in capturing fine-grained traits like coherence and argumentation, and ❸ inability to handle multimodal contexts. In the era of Multimodal Large Language Models (MLLMs), we propose EssayJudge, the first multimodal benchmark to evaluate AES capabilities across lexical-, sentence-, and discourse-level traits. By leveraging MLLMs’ strengths in trait-specific scoring and multimodal context understanding, EssayJudge aims to offer precise, context-rich evaluations without manual feature engineering, addressing longstanding AES limitations. Our experiments with 18 representative MLLMs reveal gaps in AES performance compared to human evaluation, particularly in discourse-level traits, highlighting the need for further advancements in MLLM-based AES research. The dataset and code are available at this URL.
@@
Jiamin Su1,2,*, Yibo Yan1,3,*, Fangteng Fu1, Han Zhang1, Jingheng Ye4, Xiang Liu1,3, Jiahao Huo1, Huiyu Zhou5, Xuming Hu1,2,3, [^2] Corresponding author. 1The Hong Kong University of Science and Technology (Guangzhou) 2Beijing Future Brain Education Technology Co., Ltd. 3The Hong Kong University of Science and Technology, 4Tsinghua University 5Guangxi Zhuang Autonomous Region Big Data Research Institute jsu360@connect.hkust-gz.edu.cn, yanyibo70@gmail.com, xuminghu@hkust-gz.edu.cn
@@ 2.3 Multimodal Large Language Models
MLLMs have brought significant advancements to diverse tasks and applications Xi et al. (2023); Huo et al. (2024); Yan et al. (2024d); Yan and Lee (2024); Zou et al. (2025); Dang et al. (2024). Proprietary MLLMs such as GPT-4o Hurst et al. (2024) and Gemini-1.5 DeepMind (2024b) have shown remarkable capabilities in multimodal challenges, excelling in areas such as multimodal reasoning and QA Chang et al. (2024); Yan et al. (2024c, b); Zheng et al. (2024); Yan et al. (2025a). At the same time, Open-source MLLMs have made considerable strides. For instance, LLaVA-NEXT Liu et al. (2024) utilizes a pretrained vision encoder to generate visual embeddings, which are then aligned with text embeddings through a lightweight adapter, enabling effective multimodal understanding. Similarly, MLLMs such as Qwen2-VL Wang et al. (2024), DeepSeek-VL Lu et al. (2024a), InternVL Chen et al. (2024, 2025), MiniCPM Hu et al. (2024), Ovis Lu et al. (2024b), LLaMA3 Dubey et al. (2024) and Yi-VL Young et al. (2024) implement innovative projection techniques to combine visual and textual features effectively, enabling many multimodal applications. These models showcase the growing potential of MLLMs in advancing both research and practical applications that rely on multimodal data Qu et al. (2025); Zou et al. (2024); Zhou et al. (2024); Huang et al. (2024). Therefore, we introduce EssayJudge, a novel benchmark designed to evaluate MLLMs’ capability to score essays with multimodal context, paving the way for AGI systems Xiao et al. (2024b); Tate et al. (2024); Yan et al. (2024a, 2025b).
@@ 3.1 Data Collection
3.2 Data Annotation Scheme
@@ 4.1 Experimental Setup
We meticulously categorized diverse MLLMs into distinct groups to assess their capabilities in trait-specific AES. (i) The Open-Source MLLMs category encompassed models such as Yi-VL Young et al. (2024), Qwen2-VL Wang et al. (2024), DeepSeek-VL Lu et al. (2024a), LLaVA-NEXT Liu et al. (2024), InternVL2 Chen et al. (2024), InternVL2.5 Chen et al. (2025), MiniCPM-V2.6 Hu et al. (2024), MiniCPM-LLaMA3-V2.5 Hu et al. (2024), Ovis1.6-Gemma2 Lu et al. (2024b), and LLaMA-3.2-Vision Dubey et al. (2024), each demonstrating their unique strengths and capabilities in multi-granular essay scoring. (ii) The Closed-Source MLLMs featured proprietary models like Qwen-Max Team (2024), Step-1V StepFun (2024), Gemini-1.5-Pro DeepMind (2024b), Gemini-1.5-Flash DeepMind (2024a), Claude-3.5-Haiku Anthropic (2024a), Claude-3.5-Sonnet Anthropic (2024b), GPT-4o-mini OpenAI (2024), and GPT-4o Hurst et al. (2024), providing a comparison point for the performance of models that are not publicly accessible. (iii) Lastly, the Human Performance category served as a benchmark for human-level intelligence, enabling us to assess how closely MLLMs emulate human cognitive abilities (More details in Appendix C.3). The detailed prompts for MLLMs and sources of MLLMs are provided in Appendix C.4 and C.5.
@@ 5 Conclusion
Acknowledgements
This work was supported by Open Project Program of Guangxi Key Laboratory of Digital Infrastructure (Grant Number: GXDIOP2024015); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ Appendix C Additional Experimental Details
C.1 Analysis on Scaling Law
We conducted additional experiments using a larger open-source MLLM—Qwen-VL (72B). Compared to Qwen-VL (7B), this model shows a significant performance improvement, consistent with the expected scaling law (as shown in the Table 16). However, even with increased model size, these models still underperform compared to most closed-source MLLMs, which is in line with our conclusions.
C.2 Multimodal Ablation Study
We also evaluated GPT-4o-mini. Figure 12 shows that GPT-4o-mini exhibited a decline in nine out of ten traits when image information was removed, with the exception of lexical diversity. Except for GPT-4o and GPT-4o-mini, we also evaluated text-only LLM - GPT-3.5, which also showed noticeable performance drops (as shown in Table 17 below). This finding further underscores the importance of multimodal inputs, as visual features provide essential evaluative dimensions that text-only approaches cannot capture, especially when images contain critical supporting information.
C.3 Human Performance Evaluation
@@
C.4 Prompt for MLLM Evaluation
@@ Appendix F More Multimodal Essay Scoring Examples
Table 7: Rubrics for evaluating the essay length of the essays.
@@
Table 11: Rubrics for evaluating the lexical accuracy of the essays.
@@
Table 12: Rubrics for evaluating the lexical diversity of the essays.
@@
Trait Qwen2-VL (7B) Qwen2-VL (72B)
Argument Clarity 0.17 0.34
Justifying Persuasiveness 0.10 0.52
Organizational Structure 0.14 0.48
Coherence 0.13 0.55
Essay Length 0.15 0.40
Grammatical Accuracy 0.21 0.51
Grammatical Diversity 0.16 0.55
Lexical Accuracy 0.20 0.52
Lexical Diversity 0.26 0.38
Punctuation Accuracy 0.12 0.42
Table 16: QWK values of Qwen2-VL (7B) and Qwen2-VL (72B).
Figure 12: GPT-4o-mini’s QWK values across traits for text-only and multimodal settings.
Trait GPT-4o GPT-4o-mini GPT-3.5
Argument Clarity 0.30 0.33 0.23
Justifying Persuasiveness 0.8 0.57 0.24
Organizational Structure 0.79 0.45 0.18
Coherence 0.87 0.54 0.37
Essay Length 0.70 0.46 0.35
Grammatical Accuracy 0.85 0.58 0.42
Grammatical Diversity 0.61 0.54 0.43
Lexical Accuracy 0.89 0.64 0.21
Lexical Diversity 0.89 0.56 0.24
Punctuation Accuracy 0.65 0.45 0.28
Table 17: QWK values of GPT-4o, GPT-4o-mini and GPT-3.5 without image inputs.